# MFIT : Multi-FIdelity Thermal Modeling for 2.5D and 3D Multi-Chiplet Architectures

### v2

@@ VI Conclusion
Conventional monolithic 2D chips cannot sustain the increasing performance and compute capacity demands due to increasing manufacturing costs. 2.5D and 3D multi-chiplet systems have emerged as cost-effective solutions to continue the required scaling. However, substantial compute power in a small volume intensifies the power density, leading to severe heat dissipation and thermal challenges. There is a strong need for open-source thermal modeling tools that enable researchers to analyze thermal behavior and perform thermally-aware optimizations. Re-purposing existing approaches developed for monolithic chips incurs accuracy and execution time penalties, while custom-designed singular solutions have limited scope. To fill this gap, this paper proposed MFIT, a set of multi-fidelity thermal models that span a wide range of accuracy and execution time trade-offs. Since the proposed models are consistent by construction, designers can use them throughout the design cycle, from system specification to design space exploration and runtime resource management.

### v4

@@ 6. Conclusion
Due to increasing manufacturing costs, conventional monolithic 2D chips cannot sustain the increasing performance and compute capacity demands. 2.5D and 3D chipset systems have emerged as cost-effective solutions to continue the required scaling. However, substantial compute power in a small volume intensifies the power density, leading to severe heat dissipation and thermal challenges. There is a strong need for open-source thermal modeling tools that enable researchers to analyze thermal behavior and perform thermally aware optimizations. Re-purposing existing approaches developed for monolithic chips incurs accuracy and execution time penalties, while custom-designed singular solutions have limited scope. To fill this gap, this paper proposed MFIT, a set of multi-fidelity thermal models that span a wide range of accuracy and execution time trade-offs. Since the proposed models are consistent by construction, designers can use them throughout the design cycle, from system specification to design space exploration and runtime resource management.